OpenStudy (anonymous):

5. Find the product of (x + (3+5i))2.

OpenStudy (anonymous):

please help me I do not understand this,

OpenStudy (anonymous):

I think it is x^2 + 64 ????

OpenStudy (cwrw238):

is not difficult . you just distribute 2 over whats in the brackets. in other words multiply x , + 3 and + 5i by 2 then simplify

OpenStudy (anonymous):

@cwrw238 @arabpride @LastDayWork @amriju

OpenStudy (cwrw238):

unless I've misunderstood the question - is the 2 an exponent?

OpenStudy (anonymous):

oh no, sorry its 5. (x + (3+5i))^2. squared, not times 2, sorry

OpenStudy (cwrw238):

oh ok

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

oh, and the i is an imaginary number

OpenStudy (cwrw238):

(x + (3+5i) * (x + (3+5i) = x^2 + x(3 + 5i) + x(3 + 5i) + (3 + 5i)(3 + 5i) = x^2 + 3x + 5xi + 3x + 5xi + 9 + 15i + 15i - 25 can you continue?

OpenStudy (anonymous):

so, then I just combine like terms and simplify, right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

x^2 + 6x + 10xi + 30i + 3x - 19???

OpenStudy (cwrw238):

no how did you get the -3x and -19 ?

OpenStudy (anonymous):

um ... so x^2 + 3x + 5xi + 3x + 5xi + 9 + 15i + 15i - 25 --> this is what you gave me x^2 + 10xi + 6x + 30i -19???? --> so, i saw one mistake... there is no 3x, but what makes -19 wrong? can u explain?

OpenStudy (cwrw238):

+ 9 - 25 = -16

OpenStudy (anonymous):

oh... ok, i subtracted wrong... silly me so x^2 + 10xi + 6x + 30i -16?

OpenStudy (anonymous):

Thanks soooo much!... :)

OpenStudy (anonymous):

you deserve a fan, a fan testimonial, and a medal. :D